# Latent learning: episodic memory complements parametric learning by enabling flexible reuse of experiences

### v2

@@ Abstract
When do machine learning systems fail to generalize, and what mechanisms could improve their generalization? Here, we draw inspiration from cognitive science to argue that one weakness of parametric machine learning systems is their failure to exhibit latent learning—learning information that is not relevant to the task at hand, but that might be useful in a future task. We show how this perspective links failures ranging from the reversal curse in language modeling to new findings on agent-based navigation. We then highlight how cognitive science points to episodic memory as a potential part of the solution to these issues. Correspondingly, we show that a system with an oracle retrieval mechanism can use learning experiences more flexibly to generalize better across many of these challenges. We also identify some of the essential components for effectively using retrieval, including the importance of within-example in-context learning for acquiring the ability to use information across retrieved examples. In summary, our results illustrate one possible contributor to the relative data inefficiency of current machine learning systems compared to natural intelligence, and help to understand how retrieval methods can complement parametric learning to improve generalization. We close by discussing some of the links between these findings and prior results in cognitive science and neuroscience, and the broader implications.
@@ 1 Introduction
Figure 1: Conceptual overview of the challenges of using latent information from training experiences, and how retrieval complements parametric learning to overcome them. (1(a)) The reversal curse (berglund2024reversal) is an example of how parametric learners, such as language models, consolidate information in ways that depend on the learning task and format. Models that learn a relation in one format can answer queries compatible with the learning format, but not those that reverse the relation—even though the reversed relation is latently implied by the forward one, and the models are fully capable of reversing relations to make inferences in context (lampinen2025generalization). (1(b)) Challenges of reversal are one instance of the much broader phenomenon that what is explicitly learned may also latently convey information relevant to other tasks—e.g., multi-hop reasoning, alternative goals, or answering questions in other languages. Like the reversal curse, learning on such sequences may primarily improve performance on the explicit information or goals; however, if the sequence were in context, models would readily be able to make inferences about the latent information. (1(c)) Therefore, explicit retrieval of specific experiences from nonparametric learning systems complements the broader knowledge of parametric learning—by making select, relevant experiences available in context where the latent information they contain can be used more flexibly, in ways different from the original task setting in which they were encountered. (1(d)) Thus, we will typically expect systems with solely parametric learning to perform well at new tests of knowledge that is explicit in learning experiences, but we expect selective performance advantages for episodic retrieval in tests of knowledge or tasks that are latent in learning experiences—as we demonstrate below. (These illustrative results are adapted from the simple reversals experiments below.)
@@
When do AI systems fail to generalize like natural intelligence? This question has been central to AI, and has only redoubled with the successes of deep learning over the last decade (e.g., lake2017building). For example, language models (LMs) show surprising failures to generalize to reversals of their training, e.g., failing to generalize from “X is Y’s son” in training to “Y is X’s parent” at test (berglund2024reversal)—though the models can readily make these generalizations in context (lampinen2025generalization; cf. russin2025parallel). Why do models fail to make these generalizations outside the current context, and how could this failure be addressed?
@@
Here, we draw on work in cognitive science to identify a key theme that we believe unifies and generalizes some of these past findings, and points towards a solution. Specifically, we identify the capacity for latent learning (blodgett1929effect; tolman1948cognitive) as a key gap between natural and artificial intelligence. Latent learning is the ability of a system to learn information that is not relevant to the task at hand, but that might be useful for a different future task—which can also be seen as a way of implementing a form of “prospective learning” (desilva2023prospective). We suggest that AI systems fail to exhibit latent learning in most cases—that is, AI only learns information insofar as it is relevant to the current task, and only applies information it has learned in the past insofar as it is explicitly or associatively cued by the current task. This is not to say that parametric learning does not generalize well in many cases; indeed, one of the key features of parametric learning is its ability to integrate learning over many experiences to improve generalization within a family of related tasks. However, we will argue that this type of task-driven parametric learning struggles to apply specific information from prior experiences that could be useful in a sufficiently different task at a later time. We will provide several examples and a more formal sketch of this perspective below, unifying a number of previous findings about failures of generalization and highlighting new ones.
@@
Given this argument, we again take inspiration from cognitive science to suggest a path to a solution. Specifically, in natural intelligence, episodic memory and parametric learning are complementary (mcclelland1995there; kumaran2016learning). Episodic memory seems to play an important role in some types of latent learning (kimble1968absence; myers2000latent) and other types of generalization (bayley2002medial; eichenbaum2009neurobiology). Moreover, in some of the cases where AI fails to generalize from its parametric learning, it generalizes well in context (lampinen2025generalization), suggesting that a system that uses episodic memory to reinstate the relevant information into context could generalize better. Motivated by this intuition, we test a model endowed with an oracle episodic retrieval system during both training and testing, and show that it improves on many of the latent learning failures of current methods. While at some level, this point is just a restatement of the fact that Retrieval Augmented Generation (RAG; lewis2020retrieval) can be useful, we believe that seeing it in the context of latent learning provides a new perspective on why retrieval is useful, and how current AI systems differ from natural intelligence. It therefore points to an important direction of research for AI: building more effective episodic memory and retrieval systems that approximate the role of the hippocampus in enabling generalization.
@@
  • Arguing that episodic memory (or nonparametric retrieval) can bridge this gap by bringing relevant information to more flexible online systems.
@@
  • Highlighting the importance of within-example in-context learning for learning to use information across retrieved examples.
These results shed light on some differences between natural and artificial intelligence, suggest new perspectives on extant practices like RAG, and may inspire new research in both cognitive science and AI.
@@ 2 Background
We first review some of the background and related research, and attempt to situate our work within it. To this end, we begin with the history of latent learning and complementary learning systems in cognitive science and neuroscience. We then turn to perspectives on (meta-)learning and generalization in artificial intelligence, with a particular focus on generalization within and across contexts in transformer language models.
2.1 Natural intelligence
@@
Latent learning has been suggested to depend in some cases on the medial temporal lobe (MTL; including the hippocampus) and lesions to this region impair latent learning in some cases (kimble1968absence; myers2003dissociating), as well as other types of generalization (e.g., bayley2002medial; eichenbaum2009neurobiology). (However, note that some latent learning processes may not be MTL-dependent; kimble1982further; myers2000latent.) More recent computational work has highlighted that hippocampal “preplay” might help to cache the solutions to potential future tasks (carvalho2025preemptive). Overall, these results suggest that episodic memory may contribute to some—if not all—forms of latent learning in natural intelligence.
@@
Complementary learning systems:  One strand of hippocampal research focuses on “complementary learning systems” (mcclelland1995there; kumaran2016learning)—the idea that the rapid learning and episodic memory of the hippocampus provides benefits that complement the slower, more generalized learning within the neocortex. The original argument for why the rapid-learning system is useful was that it could rapidly acquire new information without interfering with broader knowledge in the slower system; salient pieces of rapidly-acquired information could then be consolidated into the slower-learning system over time via replay.
@@
Our work suggests a potential complementary benefit of episodic memory: it could support more flexible use of past experiences in contrast to cortical learning, while cortical learning might be more closely tied to some aspects of the task or format in which the original experiences occurred. This suggestion is consistent with the lesion findings above, and more broadly compatible with arguments that the hippocampus plays a crucial role in organizing (spatial) knowledge into structures that support flexible learning and behavior (behrens2018cognitive; raju2024space)—and in linking memories together to enable generalizations such as transitive inferences (eichenbaum2009neurobiology; kumaran2012generalization) and imagination (hassabis2007patients).
@@
Transformers, context, and episodic memory in natural intelligence:  Several recent works note related connections between key computational features of transformers and natural intelligence. Most saliently, russin2025parallel highlight how capacity and flexibility tradeoffs between parametric learning and in-context learning relate to similar tradeoffs in human intelligence. More generally, various works have made computational connections between transformer attention mechanisms and episodic memory in the brain (whittington2022relating; gershman2025key). Our work complements these, by showing how these connections could point to important contributions of episodic memory to generalization, in both transformers and natural intelligence.
@@ 2.2 Artificial Intelligence
Prospective learning:  The topic of latent learning also relates to the argument of desilva2023prospective that machine learning assumes that the future will be identical in distribution to the past, but natural intelligence learns more prospectively, in a way that is oriented towards generalizing in the future. The challenges of latent learning can be interpreted as a type of prospective learning—learning information not because it is useful under the current task distribution, but because it might be useful if the task distribution changes in the future. However, the type of prospective learning focused on by De Silva et al. involves explicitly modeling how tasks might change in the future; retrieval avoids this challenge by leaving the question of which tasks may be relevant to the future (see below).
Meta-learning and in-context learning in transformers:  Another way a system can adapt to a novel future task is by learning that task once it appears; many works have explored how this capability can be meta-learned (e.g., santoro2016meta; wang2017learning), often using transformers as a memory in context (parisotto2020stabilizing; lampinen2021towards). Many works have similarly interpreted the emergent in-context learning ability in language models (brown2020language) as a type of implicit meta-learning over the training distribution (xie2022an; chan2022data; lampinen2024broader). Thus, there has been substantial interest in meta-learning (across contexts) how to learn within a single context. However, meta-learning alone does not fully address how knowledge from past contexts could be directly applied at present.
Using memory of prior contexts in agents & language models:  Several prior works in RL focused on the benefit of retrieving memory of past contexts to help adapt to present ones (e.g., ritter18been; pritzel2017neural; goyal2022retrieval; humphreys2022large). More recently, there has been substantial interest in retrieving related training documents to improve test time performance in LMs. Seminal work on Retrieval Augmented Generation (RAG; lewis2020retrieval) showed that retrieving Wikipedia articles, and adding them to the context [^1] The full RAG strategy involves additional marginalization over multiple documents, etc.; however, each document is effectively just prepended to the context., could substantially improve language model responses. More recent works have used retrieved documents to fine-tune the models (hardt2024testtime; hubotter2025efficiently) or explored incorporating retrieval into the context throughout training (e.g., khattab2020colbert; borgeaud2022improving).
@@
Patterns of generalization in language models:  Several streams of work identifying failures and successes of generalization in LMs helped inspire our investigations. First, the Reversal Curse (berglund2024reversal), and failures of multi-hop inference (balesni2024two), are intriguing failures of LMs to generalize fully the information available in their fine-tuning data. Other studies similarly find inadequate cross-linguistic transfer of knowledge in language models (qi2023cross; aggarwal2025language; goldman2025eclektic). We interpret these findings as instances of the broader phenomenon we highlight: language models learn what is explicit in the training data, in ways that may not generalize to all its latent implications for other questions (see Fig. 1(b)).
@@
3 Towards a formalization of latent learning
We next sketch a formalization of the distinction between latent learning and other types of generalization. Let $X$ be an input space, and let $X^{*}$ denote the set of finite sequences of elements from that space. Let $T^{*}$ similarly be a space of task-cue inputs. Suppose that a model is provided an input sequence $x=[x_{0},...,x_{k}]\in X^{*}$ and a task cue ($t\in T$) and is trained to both reconstruct its inputs and task cue and reproduce a task-cue-modulated mapping to an output space $f:X^{*}\times T^{*}\mapsto Y$. For example, the model might be a transformer trained via autoregressive prediction on sequences of the form $[x,t,f(x,t)]$. In a language processing setting, the input sequence $x=[x_{0},...,x_{k}]$ might be the concatenation of the individual sentences ($x_{i}$) that make up a document, the task cues $t$ might be questions about that document, and the mapping $f(x,t)$ would produce the correct answers to those questions. In an agent setting, like the latent learning experiments in rats, the inputs $x$ might consist of the identity of a maze (e.g., its name) together with the sequence of the agent’s observations and prior actions, the task cue $t$ might consist of a navigation target (e.g., a particular object in the maze), and the mapping $f(x,t)$ might consist of the optimal next action to take towards that goal.

### v3

@@ Abstract
When do machine learning systems fail to generalize, and what mechanisms could improve their generalization? Here, we draw inspiration from cognitive science to argue that one weakness of parametric machine learning systems is their failure to exhibit latent learning—learning information that is not relevant to the task at hand, but that might be useful in a future task. We show how this perspective links failures ranging from the reversal curse in language modeling to new findings on agent-based navigation. We then highlight how cognitive science points to episodic memory as a potential part of the solution to these issues. Correspondingly, we show that a system with an oracle retrieval mechanism can use learning experiences more flexibly to generalize better across many of these challenges. We also identify some of the essential components for effectively using retrieval, including the importance of within-experience in-context learning for acquiring the ability to use information across retrieved experiences. In summary, our results illustrate one possible contributor to the relative data inefficiency of current machine learning systems compared to natural intelligence, and help to understand how retrieval methods can complement parametric learning to improve generalization. We close by discussing some of the links between our work and findings in cognitive science and neuroscience—including a possible perspective on hippocampal contributions to generalization—and the broader implications.
@@ 1 Introduction
Figure 1: Conceptual overview of the challenges of using latent information from training experiences, and how retrieval complements parametric learning to overcome them. (1(a)) The reversal curse (berglund2024reversal) is an example of how parametric learners, such as language models, consolidate information in ways that depend on the learning task and format. Models that learn a relation in one format can answer queries compatible with the learning format, but not those that reverse the relation—even though the reversed relation is latently implied by the forward one, and the models are fully capable of reversing relations to make inferences in context (lampinen2025generalization). (1(b)) Challenges of reversal are one instance of the much broader phenomenon that what is explicitly learned may also latently convey information relevant to other tasks—e.g., multi-hop reasoning, alternative goals, or answering questions in other languages. Like the reversal curse, learning on such sequences may primarily improve performance on the explicit information or goals; however, if the sequence were in context, models would readily be able to make inferences about the latent information. (1(c)) Therefore, explicit retrieval of specific experiences from episodic memory complements the broader knowledge of parametric learning—by making select, relevant experiences available in context where the latent information they contain can be used more flexibly, in ways different from the original task setting in which they were encountered. (1(d)) Thus, we will typically expect systems with solely parametric learning to perform well at new tests of knowledge that is explicit in learning experiences, but we expect selective performance advantages for episodic retrieval in tests of knowledge or tasks that are latent in learning experiences—as we demonstrate below. (These illustrative results are adapted from the simple reversals experiments below.)
@@
When do AI systems fail to generalize like natural intelligence? This question has been central to AI, and has only redoubled with the successes of deep learning over the last decade (e.g., lake2017building). For example, language models (LMs) show surprising failures to generalize to reversals of their training, e.g., failing to generalize from “Plato taught Aristotle” in training to “Aristotle’s teacher was Plato’ at test (berglund2024reversal)—though the models can readily make these generalizations in context (lampinen2025generalization). Why do models fail to make these generalizations outside the current context, and how could this failure be addressed?
@@
Here, we draw on work in cognitive science to identify a key theme that we believe unifies and generalizes some of these past findings, and points towards a solution. Specifically, we identify the capacity for latent learning (blodgett1929effect; tolman1948cognitive) as a key gap between natural and artificial intelligence. Latent learning is the ability of a system to learn information that is not relevant to the task at hand, but that might be useful for a different future task—which can also be seen as a way of implementing a form of “prospective learning” (desilva2023prospective). We suggest that AI systems fail to exhibit latent learning in most cases—that is, AI only learns information insofar as it is relevant to the current task, and only applies information it has learned in the past insofar as it is explicitly cued by the current task or available through similarity-based or process-based generalization from it. This is not to say that parametric learning does not generalize well in many cases; one of the key features of parametric learning is its ability to integrate learning over many experiences to improve generalization. Indeed, there is a long history of connectionist research documenting the generalization capacities of parametric learning via semantic similarity (e.g. hinton1986learning; rogers2004semantic) or learning generalizable procedures or functions, (e.g rumelhart1986learning). However, we will argue that this type of task-driven parametric learning is distinct from the latent learning of specific information from prior experiences that could be useful in a sufficiently different task at a later time. We elaborate this distinction in the Background section below.
As a concrete example of a latent learning challenge, consider again a language model that in training learns that “Plato taught Aristotle,” but does not encounter any other facts about Plato or Aristotle. The model will likely be able to answer questions like “Who did Plato teach?” correctly with “Aristotle.” Surprisingly, however, it will not be able to answer “Who was Aristotle’s teacher?” (berglund2024reversal). Crucially though, if the forward information is available in context (e.g., by placing the training document containing the statement “Plato taught Aristotle” in the model’s context), the model will be able to answer the reverse question “Who was Aristotle’s teacher?” (lampinen2025generalization) in context. Thus, the model in some sense knows the pieces it needs to solve the reversal: it can state the forward direction or answer questions about it, and it has learned a generalizable procedure for answering reversal questions when the forward direction is available in context. However, the way that it encodes the forward direction in its parameters does not allow it to generalize to the reversal task without the forward information in context. We see this as being an instance where the forward knowledge is explicit in the data, whereas the reversal is latent—the explicitly trained information is encoded in the model’s parameters, and available in other contexts, but the model cannot use the parametric information flexibly to access what is latent in it. However, the greater flexibility of in-context learning enables accessing the latent information. We will provide a more formal outline of this perspective below, and link it to a broader set of phenomena including logical inferences and generalization to novel navigation goals—unifying a number of previous findings about successes and failures of generalization and highlighting new ones.
@@
Given this argument, we again take inspiration from cognitive science to suggest a path to a solution. Specifically, in natural intelligence, episodic memory and parametric learning are complementary (mcclelland1995there; kumaran2016learning). Episodic memory seems to play an important role in some types of latent learning (kimble1968absence; myers2000latent) and other types of generalization (bayley2002medial; eichenbaum2009neurobiology). Moreover, in some of the cases where AI fails to generalize from its parametric learning, it generalizes well in context (lampinen2025generalization), suggesting that a system that uses episodic memory to reinstate the relevant information into context could generalize better. Motivated by this intuition, we test a model endowed with an oracle episodic retrieval system during both training and testing, and show that it improves on many of the latent learning failures of current methods. Importantly, generalization from episodic retrieval does not require estimating at encoding time which information will be useful or in what way; instead, storing veridical episodes allows more flexible access at test time than can be obtained through parametric generalization alone. While at some level, this point is just a restatement of the fact that Retrieval Augmented Generation (RAG; lewis2020retrieval) can be useful, we believe that seeing it in the context of latent learning provides a new perspective on why retrieval is useful, and how current AI systems differ from natural intelligence. It therefore points to an important direction of research for AI: building more effective episodic memory and retrieval systems that approximate the role of the hippocampus in enabling generalization.
@@
  • Arguing that episodic memory (or nonparametric retrieval) can bridge this gap by making relevant information available to more flexible online systems.
@@
  • Highlighting features that contribute to learning to use retrieved information—particularly the importance of within-episode in-context learning experience for learning to use information across retrieved episodes.
  • Relating these findings to the potential role of offline and online episodic replay in enabling certain types of flexible generalization in natural intelligence.
These results shed light on some differences between natural and artificial intelligence, suggest new perspectives on extant practices such as RAG, and may inspire new research in both cognitive science and AI.
@@ 2 Background
We first review some of the background and related research, and attempt to situate our work within it. To this end, we begin with the history of latent learning and complementary learning systems in cognitive science and neuroscience. We then turn to perspectives on (meta-)learning and generalization in artificial intelligence, with a particular focus on generalization within and across contexts in transformer language models. Finally, we discuss the difference between the well-established types of similarity-based and procedural generalization in connectionist and modern models, and the latent learning challenges we discuss here.
2.1 Latent learning & episodic memory in natural intelligence
@@
Latent learning has been suggested to depend in part on the medial temporal lobe (MTL; including the hippocampus) and lesions to this region impair latent learning in some cases (kimble1968absence; myers2003dissociating), as well as other types of generalization, such as transitive inferences (eichenbaum2009neurobiology) or linguistic or semantic generalizations (bayley2002medial). (Note that some latent learning processes may not be MTL-dependent; kimble1982further; myers2000latent.) More recent work has highlighted that hippocampal “preplay” might help to cache the solutions to potential future tasks (olafsdottir2015hippocampal; carvalho2025preemptive). Overall, these results suggest that episodic memory may contribute to some—if not all—forms of latent learning in natural intelligence.
@@
Complementary learning systems:  One strand of hippocampal research focuses on “complementary learning systems” (mcclelland1995there; kumaran2016learning)—the idea that the rapid learning and episodic memory of the hippocampus provides benefits that complement the slower, more generalized learning within the neocortex. The original argument for why the rapid-learning system is useful was that it could rapidly acquire new information without interfering with broader knowledge in the slower system; these pieces of rapidly-acquired information could then be recalled when needed for immediate task performance, and integrated into the slower-learning system over time via recall or replay—which naturally results in interleaving the learning with other experiences and thereby reduce catastrophic interference.
@@
Our work suggests a potential complementary benefit of episodic memory: it could support more flexible use of past experiences in contrast to cortical learning, while cortical learning might be more closely tied to some aspects of the task or format in which the original experiences occurred. This suggestion is consistent with the lesion findings above, and more broadly compatible with arguments that the hippocampus plays a crucial role in organizing (spatial) knowledge into structures that support flexible learning and behavior (behrens2018cognitive; raju2024space; sun2023organizing)—and in linking memories together to enable generalizations such as transitive inferences (eichenbaum2009neurobiology; kumaran2012generalization) and imagination (hassabis2007patients).
@@
Transformers, context, and episodic memory in natural intelligence:  Several recent works note related connections between key features of transformers and natural intelligence. Most saliently, russin2025parallel highlight how capacity and flexibility tradeoffs between parametric learning and in-context learning relate to similar tradeoffs in human intelligence—specifically, with in-context learning having greater flexibility and potential for compositional generalization, but less capacity, relative to in-weights learning. More generally, various works have noted connections between transformer attention mechanisms and episodic memory in the brain (whittington2022relating; gershman2025key). Our work complements these, by showing how these connections could point to important contributions of episodic memory to generalization, in both transformers and natural intelligence.
@@ 2.2 Artificial Intelligence
Prospective learning:  The topic of latent learning relates to the argument of desilva2023prospective that machine learning assumes that the future will be identical in distribution to the past, but natural intelligence learns more prospectively, in a way that is oriented towards generalizing in the future. The challenges of latent learning can be interpreted as a type of prospective learning—learning information not because it is useful under the current task distribution, but because it might be useful if the task distribution changes in the future. However, the type of prospective learning focused on by De Silva et al. involves explicitly modeling how tasks might change in the future; retrieval avoids this challenge by leaving the question of which tasks may be relevant to the future (see below).
Meta-learning and in-context learning in transformers:  Another way a system can adapt to a novel future task is by learning that task once it appears; many works have explored how this capability can be meta-learned (e.g., santoro2016meta; wang2017learning), often using context in transformers as a form of memory (parisotto2020stabilizing; lampinen2021towards). Many works have similarly interpreted the emergent in-context learning ability in language models (brown2020language) as a type of implicit meta-learning over the training distribution (xie2022an; chan2022data; lampinen2024broader). Thus, there has been substantial interest in meta-learning (across contexts) to learn how to use information within a single context. However, meta-learning alone does not fully address how knowledge from past contexts could be directly applied at present.
Using memory of prior contexts in agents & language models:  Several prior works in RL focused on the benefit of retrieving memory of past contexts to help adapt to present ones (e.g., ritter18been; pritzel2017neural; goyal2022retrieval; humphreys2022large). More recently, there has been substantial interest in retrieving related training documents to improve test time performance in LMs. Seminal work on Retrieval Augmented Generation (RAG; lewis2020retrieval) showed that retrieving Wikipedia articles, and adding them to the context [^1] The full RAG strategy involves additional marginalization over multiple documents, etc.; however, each document is effectively just prepended to the context., could substantially improve language model responses. More recent works have used retrieved documents to fine-tune the models (hardt2024testtime; hubotter2025efficiently), explored incorporating retrieval into the context throughout training (e.g., khattab2020colbert; borgeaud2022improving), or allowed active retrieval via use of tools like web search (e.g. schick2023toolformer).
@@
Patterns of generalization in language models:  Several streams of work identifying failures and successes of generalization in LMs helped inspire our investigations. First, the Reversal Curse (berglund2024reversal), and failures of multi-hop inference (balesni2024two), are intriguing failures of LMs to fully exploit the information available in their fine-tuning data. Other studies similarly find inadequate cross-linguistic transfer of knowledge in language models (qi2023cross; aggarwal2025language; goldman2025eclektic). We interpret these findings as instances of the broader phenomenon we highlight: language models learn what is explicit in the training data, in ways that may not generalize to all its latent implications for other questions (see Fig. 1(b)).
@@
2.3 Similarity-based and procedural generalization from parametric learning
As noted above, we want to contrast the type of latent learning challenge that we describe here with the well-documented strengths of generalization from parametric learning. There is substantial evidence that models can and will generalize—even out of distribution—under controlled tests (e.g., abreu2025taxonomy; misra2024language). We highlight two such modes of generalization here.
Similarity-based generalization:  First, there are many studies showing a kind of similarity-based generalization from parametric learning. When entities share common features, parametric learning will build representations that reflect that similarity (hinton1986learning)—and even capture structures such as analogies on the basis of similar relations (mikolov2013distributed). These structured representations can enable some of the generalizations characteristic of human semantic cognition (rogers2004semantic), and likely underlie many of the successes of modern machine learning. We will show below that, in the presence of the right similarity cues, this type of learning can allow successful generalization in some cases that would otherwise require latent learning.
Procedural generalization or learning functions:  The second mode of generalization that parametric learning can enable is learning generalizable procedures or functions over structured inputs, that can be correctly applied in a new situation. Again, there is a long history of connectionist work studying this capability. For example, connectionist models can learn a mapping from the present tense form of a word to its past tense, in a way that generalizes to items never seen in training (rumelhart1986learning) and handles both regular and exceptional items gracefully (mcclelland2015capturing). More recently, there has also been ample evidence of stronger structural generalization from deep learning models—even in areas that have been considered challenges, such as learning generalizable procedures for variable binding (hill2020grounded), compositional generalization (lake2023human) or causal reasoning (lampinen2023passive). The flexible in-context reasoning procedures that our models demonstrate here are an instance of procedural generalization.
3 Towards a formal framework for latent learning
We next sketch a framework in which we can begin to formalize some of the ideas outlined above about the distinction between latent learning and other types of generalization. Note that this description is intended to illustrate the high-level abstractions, not to fully characterize all the forms in which a latent learning challenge may occur in practice; we discuss limitations at the end of this section.
Let $X$ be an input space, and let $X^{*}$ denote the set of finite sequences of elements from that space. Let $T^{*}$ similarly be a space of task-cue inputs. Suppose that a model is provided an input sequence $x=[x_{0},...,x_{k}]\in X^{*}$ and a task cue ($t\in T$) and is trained to both causally reconstruct its inputs and task cue and reproduce a task-cue-modulated mapping to an output space $f:X^{*}\times T^{*}\mapsto Y$. For example, the model might be a transformer trained via autoregressive prediction on sequences of the form $[x,t,f(x,t)]$. In a language processing setting, the input sequence $x=[x_{0},...,x_{k}]$ might be the concatenation of the individual sentences ($x_{i}$) that make up a document, the task cues $t$ might be questions about that document, and the mapping $f(x,t)$ would produce the correct answers to those questions. In an agent setting, like the latent learning experiments in rats, the inputs $x$ might consist of the identity of a maze (e.g., its name) together with the sequence of the agent’s observations and prior actions, the task cue $t$ might consist of a navigation target (e.g., a particular object in the maze), and the mapping $f(x,t)$ might consist of the optimal next action to take towards that goal.